OpenStudy (anonymous):

Please just answer.. Don't need a whole explanation on this. Just need to know a simple rule.

OpenStudy (anonymous):

\[(5\sqrt{3})^{2}\] I just need to know whether if the next step is just: 5 x 3 or if its: \[(5\sqrt{3})(5\sqrt{3})\]

OpenStudy (anonymous):

IT IS NOT 5x3

OpenStudy (anonymous):

IT IS THE OTHER ONE

geerky42 (geerky42):

\[(5\sqrt3)^2=(5\sqrt3)(5\sqrt3)=5^2(\sqrt3)^2\]

OpenStudy (anonymous):

ok thanks guys

OpenStudy (anonymous):

yw